OpenStudy (anonymous):

preform the indicated operation. 7/x+6/5x^2

OpenStudy (psymon):

\[\frac{ \frac{ 7 }{ (x+6) } }{ 5x ^{2} }\] This is the same as: \[\frac{ 7 }{ x+6 }\div \frac{ 5x ^{2} }{ 1 }\] It's just recognizing you can rewrite the division in this way. An entire fractional term divided by another number. Think you can solve it from there?

OpenStudy (anonymous):

is the answer 5/5x^2(x+6) ??

OpenStudy (psymon):

7 on top, otherwise correct.

OpenStudy (anonymous):

thanks :)

OpenStudy (psymon):

Yep yep